# Canonical scattering problem in topological metamaterials: Valley-Hall modes through a bend

### v1

@@ Abstract
We study the amount of backscattering of Valley Hall modes in a classical topological insulator. In reciprocal systems, the conservation of the valley index has been argued to be at the root of the high-transmission of Valley Hall modes, observed in many experimental realisations. Here, we reconsider this hypothesis by quantitatively analysing the canonical scattering problem of interface Valley Hall modes impinging on sharp bends which may or may not conserve the valley index. We consider a tight binding model of graphene ribbons with an interface and compute the reflection and transmission coefficients using a transfer matrix formalism. We find that, in all configurations, the transmission of Valley Hall modes is close to being maximal, even in cases where the valley index is not conserved. Our results reinforce the alternative interpretation of the high-transmission of Valley Hall modes in reciprocal metamaterials as a consequence of a favorable mode matching on each side of the defect and serve as a reference case for the design of Valley Hall type metamaterial.
@@ I Introduction
Topological insulators are a class of material, oftentimes constructed from an underlying periodic structure, which exhibit unique wave transport properties. First discovered in the context of electrons in solid states physics, it was then extended to the realm of classical wave: electromagnetic, elastic, acoustic and more [1, 2, 3, 4, 5]. In these systems, a non-trivial topological phase manifests itself by displaying edge waves in frequencies inside the gap of the material. A very appealing property of these edge waves is that they are able to propagate across defects or disorder without undergoing backscattering.
@@
Historically, the first example of a topological insulator was the quantum Hall effect [6]. In this system, a gap can host unidirectional edge waves, with exceptional robustness properties against defects. To obtain the necessary topology, the system must break time reversal invariance, or equivalently, break reciprocity. In condensed matter, this is ensured by an external magnetic field. Unfortunately, for classical waves, breaking reciprocity is a hard task, and always requires active materials (i.e. energy input). While the Quantum Hall effect is presumably the most studied topological insulator, another type has attracted a lot of attention since its original discovery: the Quantum Spin Hall effect [7]. The main difference is that the latter can be realized in reciprocal systems. In this case, edge waves always come in pairs moving in opposite direction. This time, backscattering immunity is achieved through the spin of the electron, which has the exotic property of inducing a phase shift of $\pi$ when time reversal operation is applied twice. This peculiarity is what can guarantee a perfect decoupling between oppositely moving modes. Unfortunately again, classical waves do not naturally possess the equivalent of the electron spin. Hence, many works in metamaterials have been devoted to realizing structures able to mimic that spin.
@@
To obtain the equivalent of an electron spin, two main approaches have been followed. The first one relies on symmetries to obtain degenerate energy levels, acting as pseudo-spins [8, 9, 10]. The second one on the contrary, starts with a configuration displaying several inequivalent Dirac points and break inversion symmetry [11, 12, 13, 14]. That second case, which is the focus of this work, is called the Valley Hall effect. The reason is that the symmetry breaking turns the Dirac points into small gaps with valley shaped dispersion relations. Each valley carries a topological charge, and is essentially decoupled from the others. A typical configuration to illustrate this decoupling is the scattering of valley edge modes on sharp bends and various possible combinations, such as $Z$ and $\Omega$ shaped paths. Many works have observed a very high transmission across these bends, which seems to originate from valley conservation[15, 16, 17, 18, 19, 20].
@@
In the last years, it was established in a couple of works that backscattering immunity is never exact in these time reversal invariant topological insulators with classical waves [21, 22, 23, 24, 25, 26]. For instance, in the Valley Hall effect decoupling between valleys is always approximate. Hence, the question was raised whether the observed high transmission across bends is due to a topological protection or more incidentally, a good mode matching on each side of the bend. To settle this question, quantitative analysis of the scattering across bands must be conducted. In this work, we analyse this scattering problem in the context of a tight binding model of a Valley Hall system.
@@
The motivation of this work is to see this specific model as a canonical scattering problem, which could serve as a reference case for Valley Hall type metamaterials[27] and to further understand the necessary conditions for high-transmission across sharp bends [28, 29, 30]. Here we investigate the propagation of Valley Hall states in graphene ribbons with two possible interfaces and quantify the amount of backreaction when Valley Hall states scatter through sharp turns with an angle of $\pi/3$ and $32\pi/$ 3. Reflection and transmission coefficients are defined via a scattering problem and computed numerically up to machine-precision using the transfer matrix formalism. We show that Valley Hall states are well transmitted through the two possible turns, independently of the angle. Our results support the mode-matching interpretation of the observed high-transmission rather than the topological argument of valley conservation.
@@
The paper is organised as follows: In Sec. II we discuss the valley Hall effect and its realisation in a lattice model. We describe the band structure and the properties of interface modes in graphene ribbons with an interface. In Sec. III, we review the transfer matrix formalism which allows us to quantify the amount of backscattering. In Sec. IV we present the quantitative analysis of the scattering of valley Hall states through a turn at different angles. Sec. V shows an extension of our formalism to quantitatively analyse other lattice ribbons with an application to graphene ribbons with armchair edges. We finish our study by a discussion of our results in Sec. VI.
@@ II Lattice Model
In the following, we will investigate the realisation and scattering of valley Hall modes in a discrete honeycomb lattice system. In this model, the system is described by a set of complex amplitudes associated with each node of the lattice and coupled to their nearest neighbours. Such approach is commonly used in condensed matter as it describes the tight-binding approximation of the propagation of electrons in crystals. Discrete models play an important role in modelling continuous classical system in asymptotic regimes such as low-frequency phononics and photonics crystals [31], mechanical lattices [32] or in the mono-mode approximation of equal-lengths acoustic networks [33, 34, 35].
@@ II.1.2 Opening up the gap
In order to realise the valley Hall effect, we now break a spatial symmetry of the system by introducing an onsite potential $u$, which takes opposite values on the A and B sites. This addition breaks inversion symmetry of the lattice which is now left invariant under $32\pi/$ 3 rotation, reflections along axes passing through vertices of the lattice and the translations along the vectors $\bm{v}_{1}$ and $\bm{v}_{2}$. The lattice with broken spatial symmetry is shown in Fig. 2-a). We choose the black sites to have an onsite potential $+u$ and the white sites to have an onsite potential $-u$. In general $u$ will be chosen positive unless specified.
@@
The energy eigenvalues are given by $E_{\pm}=\pm\sqrt{u^{2}+|f|^{2}}$, which leads to a band structure presenting a gap of magnitude $2u$, as shown in Fig. 2-c) & d). While the energy eigenvalues are identical in the two valleys $K$ and $K^{\prime}$, the modes in each valley are distinct. This becomes apparent when one computes the Berry curvature, shown in Fig. 2-b) (see Appendix A for definition and computational details). As one can see, the Berry curvature has opposite signs in the two valleys and it can be used as a pseudo-spin. This construction together with arguments on the conservation of the valley index is at the basis of valleytronics and of the topological protection of valley Hall modes, which we shall investigate quantitatively in graphene ribbons in the next sections.
@@ II.2.1 Graphene ribbons
When considering graphene ribbons, one has to decide how to “cut” the ribbons within the periodic graphene lattice. In other words, one has to choose the shape of the ribbon edges. There are infinitely many ways to cut the graphene lattice and authors have studied ribbons with various edges. Amongst all the possibilities stem three canonical choices: zigzag edges, bearded edges and armchair edges, which arise as straight cuts along high-symmetry directions of the honeycomb lattice [36]. Zigzag and bearded edges are obtained when cutting along the direction of a lattice vector while the armchair is obtained when cutting in a direction perpendicular to a lattice vector. Since the main focus of our work concerns modes localised near the interface and away from the edges, the direction of the ribbon interface is much more relevant that the edge type hence we will focus in the following on zigzag edges. The case of armchair ribbons is discussed in Sec. V.
@@ II.2.2 Graphene ribbons with onsite potential
Figure 4: Band structure and edge modes of graphene ribbons with zigzag edges and onsite potential $u=0.5$. Panel (a) shows the dispersion relation which exhibits a band gap. Panel (b) shows an edge mode located near the upper edge of the ribbon. The area of the dots represents the amplitude of the mode while the color illustrates the phase.
@@ II.2.3 Graphene ribbons with an interface
This separation becomes apparent at the edges of the Brillouin zone, where we can distinguish a set of special energy values depending on the type of interface considered. Those energy values can be extracted from the limit $s(k)=0$, corresponding to the fully dimmerized Rice-Mele chain, and calculations are presented in Appendix B. When $k\rightarrow\pm\pi$ and therefore $s(k)\rightarrow 0$, we can see that the bulk modes accumulate around the energies $E=\pm\sqrt{1+u^{2}}$. For a bridge interface the edges modes have energy $E=-u$ and the interface modes have energies $E=u\pm 1$. For a zigzag interface the edges modes have energy $E=u$ and the interface modes have energies $E=\sqrt{1+u^{2}}$ and join the bulk modes.
@@
Figs. 5-c) and 6-c) show interface states in both configurations. Note that the addition of an interface restores the mirror symmetry of the Rice-Mele chain associated with the ribbon supercell. The eigenmodes of the chain (and therefore of the ribbon provided the phase transformation given in Eqs. (16) & (17)) can again be decomposed into a symmetric and antisymmetric sector. It can be shown that the interface mode in the band gap is an antisymmetric mode. The energy and symmetry properties of the interface states can be understood in the limit of ribbons with infinite width which is presented in Appendix B.
@@
where $\bm{U}_{A/B}$ are diagonal matrices containing the values of the onsite potential on the A/B sites. These equation can be rewritten as
@@ IV.1 Geometrical configurations
Before discussing the scattering properties of the various configurations, we begin by noticing geometrical properties of the different scenarios.
@@
[…]
  • The onsite potential at the interface changes sign on each side of the bend in the case of a $\pi/3$ bend while it remains unchanged in the case of a $32\pi/$ 3 bend.
@@
Figure 9: Different geometrical configurations combining the two possible interfaces and the two possible angles for the bend. The solid red line represents the interface: zigzag in Panels a) and b) and bridge in Panels c) and d). The green dashed line represents the axis involved in the reflection symmetries of the various configurations.
@@ IV.2.1 Spectrum
We recall from the last part of Sec. II that the chiral operator, $\bm{\Gamma}$, acts as
$\bm{\Gamma}\psi^{\pm}_{L/R}(E,u)=\psi^{\pm}_{L/R}(-E,-u),$ (68)
@@
$\bm{\Gamma}\Psi_{n}(E,u)$ $\underset{n\rightarrow-\infty}{\sim}$ $\psi^{+}_{L}(-E,-u)+R(E,u)\ \psi^{-}_{L}(-E,-u),$ (69)
@@
Applying the mirror operator, $\bm{\mathcal{M}}$, which acts on the eigenmodes as
@@
[…]
@@ IV.2.2 Visualisation of scattering solutions
Fig. 13 shows the reconstructed scattering solution for a bridge interface and a bend with an angle of $\pi/3$. Note that the sites closest to the interface on the ribbon to the left of the bend have a positive onsite potential while the ones on the right of the bend have a negative onsite. This implies in particular that the interface states are anti-symmetric in the left ribbon and symmetric in the right ribbon, as shown in Appendix B. This transition between anti-symmetric to symmetric modes is clearly visible in the reconstruction.

### v2

@@ Abstract
We study the amount of backscattering of Valley Hall modes in a classical topological insulator. In reciprocal systems, the conservation of the valley index has been argued to be at the root of the high-transmission of Valley Hall modes, observed in many experimental realisations. Here, we reconsider this hypothesis by quantitatively analysing the canonical scattering problem of interface Valley Hall modes impinging on sharp bends which may or may not conserve the valley index. We consider a tight binding model of graphene ribbons with an interface and compute the reflection and transmission coefficients using a transfer matrix formalism. We find that, in all configurations considered, the transmission of Valley Hall modes is close to being maximal, even in cases where the valley index is not conserved. Hence there appears to be no correlation between valley conservation and good transmission. Our results serve as a reference case for the design of Valley Hall type metamaterial.
@@ I Introduction
Topological insulators are a class of material, oftentimes constructed from an underlying periodic structure, which exhibit unique wave transport properties. First discovered in the context of electrons in solid states physics, it was then extended to the realm of classical wave: electromagnetic, elastic, acoustic and more [1, 2, 3, 4, 5, 6, 7]. In these systems, a non-trivial topological phase manifests itself by displaying edge waves in frequencies inside the gap of the material. A very appealing property of these edge waves is that they appear to be able to propagate across defects or disorder without undergoing backscattering.
@@
Historically, the first example of a topological insulator was the quantum Hall effect [8]. In this system, a gap can host unidirectional edge waves, with exceptional robustness properties against defects. To obtain the necessary topology, the system must break time reversal invariance, or equivalently, break reciprocity. In condensed matter, this is ensured by an external magnetic field. Unfortunately, for classical waves, breaking reciprocity is a hard task, and always requires active materials (i.e. energy input). While the Quantum Hall effect is presumably the most studied topological insulator, another type has attracted a lot of attention since its original discovery: the Quantum Spin Hall (QSH) effect [9]. The main difference is that the latter can be realized in reciprocal systems. In this case, edge waves always come in pairs moving in opposite direction. This time, backscattering immunity is achieved through the spin of the electron, which has the exotic property of inducing a phase shift of $\pi$ when time reversal operation is applied twice. This peculiarity is what can guarantee a perfect decoupling between oppositely moving modes [10]. Unfortunately again, classical waves do not naturally possess the equivalent of the electron spin. Hence, many works in metamaterials have been devoted to realizing structures able to mimic that spin with the hope that they would protect their associated topological edge waves from backscattering.
@@
To obtain the equivalent of an electron spin, two main approaches have been followed. The first one relies on symmetries to obtain degenerate energy levels, acting as pseudo-spins [11, 12, 13]. This historical analogue QSH effect has been extensively studied theoretically and experimentally in mechanical systems [14], acoustics [15, 12] and photonics [16, 17] and it has been established that QSH modes exhibit a reduced backscattering against various type of defects. The second one on the contrary, starts with a configuration displaying several inequivalent Dirac points and break inversion symmetry [18, 19, 20, 21]. That second case, which is the focus of this work, is called the Valley Hall effect. The reason is that the symmetry breaking turns the Dirac points into small gaps with valley shaped dispersion relations. Each valley carries a topological charge, and is essentially decoupled from the others. A typical configuration to illustrate this decoupling is the scattering of valley edge modes on sharp bends and various possible combinations, such as $Z$ and $\Omega$ shaped paths. Similarly to the case of QSH devices, many works have observed a very high transmission of valley Hall modes across these bends, which seems to originate from valley conservation[22, 23, 24, 25, 26, 27, 28]. While both the analogue QSH and VH effects aim at creating pseudo-spin to protect the topological modes, they do so in fundamentally different ways. The QSH designs are based on a 6-fold rotational symmetry while VH systems break inversion symmetry. The two approaches have been compared in a few specific settings, and their robustness to backscattering seems to depend substantially on the type of defects and their implementations [29, 30]. This observation indicates that the scattering properties are not solely dependent on the topological properties of the modes but also on the symmetries of the underlying systems.
@@
Along these lines and in the last years, it was established in a couple of works that backscattering immunity is never exact in these time reversal invariant topological insulators with classical waves [29, 30, 31, 32, 33, 34]. For instance, in the Valley Hall effect decoupling between valleys is always approximate, while, in the QSH effect, the introduction of a defect necessarily breaks the spatial symmetries needed for the construction of the QSH topological invariant. Hence, the question was raised whether the observed high transmission across bends is due to a topological protection or to an accidental (i.e. not protected by symmetry) good mode matching on each side of the bend. Such question is particularly relevant, since there exist scattering configurations that inevitably require the non-conservation of the valley index, as we will see below. To settle this question, quantitative analysis of the scattering across bands must be conducted. In this work, we analyse this scattering problem in the context of a tight binding model of a Valley Hall system. Such systems can be seen as a proxy for a photonic or phononic metamaterial but are also pertinent for condensed matter systems.
@@
The motivation of this work is to see this specific model as a canonical scattering problem, which could serve as a reference case for Valley Hall type metamaterials[35] and to further understand the necessary conditions for high-transmission across sharp bends [36, 37, 38]. Here we investigate the propagation of Valley Hall states in graphene ribbons with two possible interfaces. Using a transfer matrix formalism, we quantify the amount of backscattering when Valley Hall states scatter through sharp turns with an angle of $\pi/3$ and $32\pi/$ 3. The main difference between the two configurations is that it is possible to conserve the valley index for the $32\pi/$ 3 turn while it is not for the $\pi/3$ case. We show that Valley Hall states are well transmitted through the two possible turns, independently of the angle. Our results support the idea that valley conservation is not correlated to a high transmission across turns.
@@
The paper is organised as follows: In Sec. II we discuss the valley Hall effect and its realisation in a lattice model. We describe the band structure and the properties of interface modes in graphene ribbons with an interface. In Sec. III, we review the transfer matrix formalism which allows us to quantify the amount of backscattering. In Sec. IV we present the quantitative analysis of the scattering of valley Hall states through a turn at different angles. We finish our study by a discussion of our results in Sec. V. The main text is complemented by supplementary materials which contain information about technical details and an extension of our formalism to quantitatively analyse other lattice ribbons with an application to graphene ribbons with armchair edges.
@@ II Lattice Model
In the following, we will investigate the realisation and scattering of valley Hall modes in a discrete honeycomb lattice system. In this model, the system is described by a set of complex amplitudes associated with each node of the lattice and coupled to their nearest neighbours. Such approach is commonly used in condensed matter as it describes the tight-binding approximation of the propagation of electrons in crystals. Discrete models play an important role in modelling continuous classical system in asymptotic regimes such as low-frequency phononics and photonics crystals [39], mechanical lattices [40] or in the mono-mode approximation of equal-lengths acoustic networks [41, 42, 43]. It is worth noting that tight-binding models, usually arising in a regime of weak coupling in condensed matter, have found applications in systems with long range couplings such as elastic systems which exhibit a very strong coupling between different resonators as well as between various kind of elastic waves (transverse and longitudinal) [44, 45].
The building blocks of discrete models are : 1) a set of sites where the relevant field is located and has a (complex) amplitude $A_{n}$ (where $n\in\mathbb{N}$ denotes the various sites), and 2) hopping coefficients which characterise the couplings between the different sites of the lattice. Those two quantities, the field amplitudes and the hopping coefficients, are quite generic but they can be related to physical quantities depending on the system of interest. For example, in the context of acoustic waves, the discrete model can be seen as an acoustic network of tubes where the field amplitude can be understood as the pressure field at the nodes of the network and the hopping coefficients are related to the different cross section areas of the tubes in the network [41]. In spring-mass networks, the amplitude is related to the displacements of the masses while the hoppings are linked to the spring constants and the masses [46], while in photonic crystals they can be related to the electric field amplitude and the dielectric constants of the system [39].
@@ II.1.2 Opening up the gap
In order to realise the valley Hall effect, we now break a spatial symmetry of the system by introducing an onsite potential $u$, which takes opposite values on the A and B sites. This addition breaks inversion symmetry of the lattice which is now left invariant under $32\pi/$ 3 rotation, reflections along axes passing through vertices of the lattice and the translations along the vectors $\bm{a}$ and $\bm{b}$. The lattice with broken spatial symmetry is shown in Fig. 2-a). We choose the black sites to have an onsite potential $+u$ and the white sites to have an onsite potential $-u$. In general $u$ will be chosen positive unless specified.
@@
The energy eigenvalues are given by $E_{\pm}=\pm\sqrt{u^{2}+|f|^{2}}$, which leads to a band structure presenting a gap of magnitude $2u$, as shown in Fig. 2-c) & d). While the energy eigenvalues are identical in the two valleys $K$ and $K^{\prime}$, the modes in each valley are distinct. This becomes apparent when one computes the Berry curvature, shown in Fig. 2-b) (see supplemental material Sec. A for definition and computational details). As one can see, the Berry curvature has opposite signs in the two valleys and it can be used as a pseudo-spin. This construction together with arguments on the conservation of the valley index is at the basis of valleytronics and of the topological protection of valley Hall modes, which we shall investigate quantitatively in graphene ribbons in the next sections.
@@ II.2.1 Graphene ribbons
When considering graphene ribbons, one has to decide how to “cut” the ribbons within the periodic graphene lattice. In other words, one has to choose the shape of the ribbon edges. There are infinitely many ways to cut the graphene lattice and authors have studied ribbons with various edges. Amongst all the possibilities stem three canonical choices: zigzag edges, bearded edges and armchair edges, which arise as straight cuts along high-symmetry directions of the honeycomb lattice [47]. Zigzag and bearded edges are obtained when cutting along the direction of a lattice vector while the armchair is obtained when cutting in a direction perpendicular to a lattice vector. Since the main focus of our work concerns modes localised near the interface and away from the edges, the direction of the ribbon interface is much more relevant that the edge type hence we will focus in the following on zigzag edges. The case of armchair ribbons is discussed in the supplemental material Sec. E.
@@ II.2.2 Graphene ribbons with onsite potential
Figure 4: Band structure and edge modes of graphene ribbons with zigzag edges and onsite potential $u=0.5$. Panel (a) shows the dispersion relation which exhibits a band gap. Panel (b) shows two edge modes located either on the lower or the upper edge of the ribbon. The area of the dots represents the amplitude of the mode while the color illustrates the phase.
@@ II.2.3 Graphene ribbons with an interface
This separation becomes apparent at the edges of the Brillouin zone, where we can distinguish a set of special energy values depending on the type of interface considered. Those energy values can be extracted from the limit $s(k)=0$, corresponding to the fully dimmerized Rice-Mele chain, and calculations are presented in the supplemental material Sec. B. When $k\rightarrow\pm\pi$ and therefore $s(k)\rightarrow 0$, we can see that the bulk modes accumulate around the energies $E=\pm\sqrt{1+u^{2}}$. For a bridge interface the edges modes have energy $E=-u$ and the interface modes have energies $E=u\pm 1$. For a zigzag interface the edges modes have energy $E=u$ and the interface modes have energies $E=\sqrt{1+u^{2}}$ and join the bulk modes.
@@
Figs. 5-c) and 6-c) show interface states in both configurations. Note that the addition of an interface restores the mirror symmetry of the Rice-Mele chain associated with the ribbon supercell. The eigenmodes of the chain (and therefore of the ribbon provided the phase transformation given in Eqs. (16) & (17)) can again be decomposed into a symmetric and antisymmetric sector. It can be shown that the interface mode in the band gap is an antisymmetric mode. The energy and symmetry properties of the interface states can be understood in the limit of ribbons with infinite width which is presented in Sec. B of the supplemental material.
@@
where $U_{A/B}$ are diagonal matrices containing the values of the onsite potential on the A/B sites. These equation can be rewritten as
@@ IV.1 Geometrical configurations
Before discussing the scattering properties of the various configurations, we begin by noticing general properties of the different scenarios.
@@
  • The $\pi/3$ angle configurations are symmetric under the operation $\mathcal{M}\Gamma$, where $\mathcal{M}$ is the operator corresponding to mirror symmetry with respect to the green dashed line in Figs. 9-a) and c), and $\Gamma$ is the chiral operator of Eq. (18). Indeed, we notice that the mirror operation flips the role of A and B, which means a flip of sign of the on-site potential. Subsequently applying $\Gamma$ flips back the on-site potential, so the overall configuration is invariant under the action of $\mathcal{M}\Gamma$.
  • The onsite potential at the interface changes sign on each side of the bend in the case of a $\pi/3$ bend while it remains unchanged in the case of a $32\pi/$ 3 bend. This implies in particular that the scattering of a valley Hall mode in this configuration will necessarily flip its valley index since the reflected and transmitted waves lie in the opposite valley, as represented by the arrows in Fig. 9.
@@
Figure 9: Different geometrical configurations combining the two possible interfaces and the two possible angles for the bend. The solid red line represents the interface: zigzag in Panels a) and b) and bridge in Panels c) and d). The green dashed line represents the axis involved in the reflection symmetries of the various configurations. The arrows represent the direction of propagation of the interface modes and their color indicates the valley index (see Fig. 1).
@@ IV.2.1 Spectrum
We recall from the last part of Sec. II that the chiral operator, $\Gamma$, acts as
$\Gamma\psi^{\pm}_{L/R}(E,u)=\psi^{\pm}_{L/R}(-E,-u),$ (68)
@@
$\Gamma\Psi_{n}(E,u)$ $\underset{n\rightarrow-\infty}{\sim}$ $\psi^{+}_{L}(-E,-u)+R(E,u)\ \psi^{-}_{L}(-E,-u),$ (69)
@@
Applying the mirror operator, $M$, which acts on the eigenmodes as
@@
Hence from the unicity of the scattering solution, $M\Gamma\Psi_{n}(E,u)$ is a solution to the scattering at energy $-E$, with reflection and transmission coefficients $R(-E,u)$ and $T(-E,u)$. We can therefore read the following relations:
@@ IV.2.2 Visualisation of scattering solutions
Fig. 13 shows the reconstructed scattering solution for a bridge interface and a bend with an angle of $\pi/3$. Note that the sites closest to the interface on the ribbon to the left of the bend have a positive onsite potential while the ones on the right of the bend have a negative onsite. This implies in particular that the interface states are anti-symmetric in the left ribbon and symmetric in the right ribbon, as shown in Sec. B of the supplemental material. This transition between anti-symmetric to symmetric modes is clearly visible in the reconstruction.